# Assessing and Learning Alignment of Unimodal Vision and Language Models

### v1

@@ Abstract
How well are unimodal vision and language models aligned? Although prior work have approached answering this question, their assessment methods do not directly translate to how these models are used in practical vision-language tasks. In this paper, we propose a direct assessment method, inspired by linear probing, to assess vision-language alignment. We identify that the degree of alignment of the SSL vision models depends on their SSL training objective, and we find that the clustering quality of SSL representations has a stronger impact on alignment performance than their linear separability. Next, we introduce Swift Alignment of Image and Language (SAIL), a efficient transfer learning framework that aligns pretrained unimodal vision and language models for downstream vision-language tasks. Since SAIL leverages the strengths of pretrained unimodal models, it requires significantly fewer ($\sim$6%) paired image-text data for the multimodal alignment compared to models like CLIP which are trained from scratch. SAIL training only requires a single A100 GPU, $\sim$5 hours of training and can accommodate a batch size up to 32,768. SAIL achieves 73.4% zero-shot accuracy on ImageNet (vs. CLIP’s 72.7%) and excels in zero-shot retrieval, complex reasoning, and semantic segmentation. Additionally, SAIL improves the language-compatibility of vision encoders that in turn enhance the performance of multimodal large language models. The entire codebase and model weights are open-source: Project Page.
@@ 1 Introduction
Figure 1: Conceptual Overview: Alignment probing evaluates the alignment potential of two pretrained uni-modal models. Akin to linear probing, only the linear alignment layers are trained while the backbone models remain frozen.
The integration of language and vision is pivotal in advancing models for zero-shot open-vocabulary computer vision tasks [27, 28, 20, 21]. This raises a key question: “To what extent can unimodal visual and language models be aligned with each other?” In particular, how do unimodal representations impact cross-modal alignment: Do larger unimodal models trained on extensive datasets yield better cross-modal alignment? Does the choice of self-supervised learning (SSL) method play a critical role in determining the alignment strength? What property of SSL representation correlates the most with cross-modal alignment performance: is it linear separability or the clustering quality?
Although prior work have approached answering the question of “To what extent are pretrained unimodal models aligned with each other”, their assessment methods do not directly translate to how these models are used in practical vision-language tasks. For instance, Huh et al. [19] assess cross-modal alignment using mutual nearest-neighbor metrics, indicating a certain level of alignment across models trained on separate modalities. However, this assessment method serves only as a proxy method as it focuses on relative ordering within each modality rather than directly measuring cross-modal distances. The latter is precisely how inference with vision-language models (VLMs) is conducted, so we believe directly measuring cross-modal distances is a more direct measurement of alignment performance.
@@
2 Assessing Alignment between Unimodal Models
@@ 2.2 Results and Findings
We further investigate which properties of SSL representations best support alignment. There are two standard metrics to probe the quality of SSL representations [16, 6, 4, 33]: 1) k-NN classifier, which measures non-linear separability and clustering quality of the SSL representations, i.e., how well representations of the same concept cluster together, 2) Linear Probing which measures linear separability, i.e., how effectively the features can be separated by a linear boundary. Our analysis reveals a strong linear correlation between k-NN performance and alignment score, as illustrated by an approximated line in the figure. In contrast, Linear Probing (Appendix) shows a weaker correlation with alignment scores . Specifically, Pearson’s correlation between alignment accuracy (computed using our proposed metric) and ImageNet classification accuracy is 0.991 for k-NN classification and 0.847 for linear probing, highlighting that non-linear separability (i.e., clustering quality) matters more than linear-separability for image-text alignment.
@@
3 Learning Alignment between Unimodal Models
@@ 3.2.2 Zero-Shot Image-Text Retrieval
Cross-modal retrieval is more challenging than image recognition, as it requires complex scene understanding, including spatial relationships, context, background, activities, and more. Unlike models like CLIP that lack complex reasoning capabilities [41] as they train from scratch using large, noisy image-text datasets with short captions mentioning just object names, SAIL leverages a robust pretrained language encoder, trained for complex language understanding. With minimal alignment learning on a significantly smaller image-text dataset, SAIL achieves significant cross-modal understanding improvements.
@@
We further analyze the image-image cosine similarity on the MMVP benchmark [43], which consists of 150 image pairs. These pairs are selected to test subtle differences in orientation, perspective, quantity, color, and contextual details. Previous findings [43] indicate that CLIP struggles with such distinctions, often assigning very high similarity scores even when condition varies, whereas DINOv2 effectively captures these nuances. As shown in Fig. 6, we found that SAIL’s cosine similarity distribution closely matches DINOv2’s, suggesting it retains DINOv2’s fine-grained visual acuity. By combining DINOv2’s visual precision with NV2’s linguistic depth, SAIL proves to be a powerful vision-language foundation model for nuanced visual discrimination.
@@ 3.2.3 Open-Vocabulary Semantic Segmentation
[…]
@@
[…]
@@ 3.2.4 Language-compatible Visual Representation
Figure 9: CLIP and LiT require training full models, making them resource-intensive. LiT and ShareLock freeze vision models during training, unable to yield language-compatible visual features crucial for MLLMs. Lastly, CLIP and LiT achieves modality-specific representations and cross-modal alignment simultaneously, incapable for alignment evaluation. Proposed SAIL meets all three requirements.
@@ 4 Related Work
Vision-Language Models
Foundational VLMs like CLIP are widely used across vision-language tasks such as retrieval, classification, and segmentation in zero-shot settings, and they serve as essential components in multimodal generative models. For instance, CLIP functions flexibly as a text encoder in text-to-image generation [36, 35] and as a vision encoder in MLLMs [27, 9, 42], excelling through its aligned multimodal representations.
Alignment between unimodal models
Recent studies reveal that alignment can emerge within unimodal models even without explicitly aligning them with each other. Huh et al. [19] use mutual nearest-neighbor metrics to suggest that models across modalities align to a shared statistical reality. Maniparambil et al. [30] find that vision encoders exhibit high semantic similarity with language encoders using Centered Kernel Distance. While these studies imply inherent alignment within unimodal models, they rely on proxy measurements without directly assessing cross-modal distance for individual image-text pairs. Concurrent work by Sharma et al. [37] maps language models to visual representation spaces, identifying large-scale decoder-based LLMs as ideal candidates for vision-centric text representation. Our approach directly measures the cross-modal distance for individual image-text pairs to quantitatively examine how modality-specific features impact alignment.
@@
Training VLMs from scratch requires extensive datasets and computational resources, especially for contrastive models like CLIP, which demand large batch sizes, and MLLMs such as Gemini [40] and Fuyu [2] that involve training LLMs. Previous studies have shown that strong pre-trained vision models and LLMs can be efficiently aligned using linear layers [31, 27, 1], resulting in powerful MLLMs.
On the other hand, efficient training of foundational VLMs like CLIP remains underexplored. One line of work reduces data requirements by using improved captions from LLMs [12] or MLLMs [48], though computational demands remain high. Other methods combine pre-trained unimodal models to reduce data and compute needs. For instance, LiT [45] aligns a frozen vision model with a language model trained from scratch, which still requires substantial pretraining. ShareLock [37] further introduces a tunable MLP layer over the frozen language model to achieve alignment with fewer parameters. However, these methods have limited alignment performance and do not enhance the base vision encoder, restricting the ability to transfer improved vision-language alignment capabilities to MLLMs. As shown in Fig. 9, SAIL stands out by meeting three key points compared to other frameworks: efficient training, learning language-compatible visual features, and a more direct alignment evaluation method consistent with how inference is done with such vision-language models.
@@ 5 Conclusion
This work proposes a alignment probing framework, inspired by linear probing, to evaluate cross-modal alignment between pretrained unimodal vision and language models and to explore how modality-specific features impact this alignment. Our results show that the clustering quality of self-supervised learning features, assessed by the kNN classifier, is crucial for effective alignment. Additionally, high-performing language models are found to be essential for complex reasoning in vision-language tasks.
Building on these findings, we introduce the SAIL framework, which achieves optimal vision-language alignment with minimal human annotation. By utilizing pretrained SSL models, fewer resources, and a streamlined training setup, SAIL excels in zero-shot classification, cross-modal retrieval, complex visio-linguistic reasoning, and open-vocabulary segmentation, surpassing large-scale models like CLIP trained on 400M image-text data. It also enables learning of visual encoders that are more compatible to be used in multimodal language models.
This study underscores the potential of efficient alignment strategies to advance practical vision-language integration, emphasizing the advantages of pretrained unimodal models and minimal human supervision. We hope our findings will accelerate foundational research on vision-language models focused on topics such as architecture, losses, and data, particularly for academic groups with limited resources.
@@ Appendix A Reproducibility Statement
To ensure the reproducibility of our work, we are committed to making all training code, datasets, and model weights publicly available. Detailed documentation will accompany the codebase to facilitate easy replication of our experiments. Hyperparameter settings, training configurations, and any preprocessing steps will also be thoroughly outlined. By providing these resources, we aim to promote transparency, enable future research, and support the broader community in building upon our work.

### v2

@@ Abstract
How well are unimodal vision and language models aligned? Although prior work have approached answering this question, their assessment methods do not directly translate to how these models are used in practical vision-language tasks. In this paper, we propose a direct assessment method, inspired by linear probing, to assess vision-language alignment. We identify that the degree of alignment of the SSL vision models depends on their SSL training objective, and we find that the clustering quality of SSL representations has a stronger impact on alignment performance than their linear separability. Next, we introduce Swift Alignment of Image and Language (SAIL), a efficient transfer learning framework that aligns pretrained unimodal vision and language models for downstream vision-language tasks. Since SAIL leverages the strengths of pretrained unimodal models, it requires significantly fewer ($\sim$6%) paired image-text data for the multimodal alignment compared to models like CLIP which are trained from scratch. SAIL training only requires a single A100 GPU, $\sim$5 hours of training and can accommodate a batch size up to 32,768. SAIL achieves 73.4% zero-shot accuracy on ImageNet (vs. CLIP’s 72.7%) and excels in zero-shot retrieval, complex reasoning, and semantic segmentation. Additionally, SAIL improves the language-compatibility of vision encoders that in turn enhance the performance of multimodal large language models. [^1] Project page: https://lezhang7.github.io/sail.github.io/
@@ 1 Introduction
Figure 1: Conceptual Overview: Alignment probing evaluates the alignment potential of two pretrained uni-modal models.
The integration of language and vision is pivotal in advancing models for computer vision tasks [27, 28, 19, 20, 46]. This raises a key question: “To what extent can unimodal visual and language models be aligned with each other?" In particular, how do unimodal representations impact cross-modal alignment: Do larger unimodal models trained on extensive datasets yield better cross-modal alignment? Does the choice of self-supervised learning (SSL) method play a critical role in determining the alignment strength? What property of SSL representation correlates the most with cross-modal alignment performance: is it linear separability or the clustering quality?
Although prior work have approached answering the question of “To what extent are pretrained unimodal models aligned with each other", their assessment methods do not directly translate to how these models are used in practical vision-language tasks. For instance, Huh et al. [17] assess cross-modal alignment using mutual nearest-neighbor metrics, indicating a certain level of alignment across models trained on separate modalities. However, this assessment method serves only as a proxy method as it focuses on relative ordering within each modality rather than directly measuring cross-modal distances. The latter is precisely how inference with vision-language models (VLMs) is conducted, so we believe directly measuring cross-modal distances is a more direct measurement of alignment performance.
@@
2 Part I: Assessing Alignment between Unimodal Models
@@ 2.2 Results and Findings
We further investigate which properties of SSL representations best support alignment. There are two standard metrics to probe the quality of SSL representations [14, 5, 3, 34]: 1) k-NN classifier, which measures non-linear separability and clustering quality of the SSL representations, i.e., how well representations of the same concept cluster together, 2) Linear Probing which measures linear separability, i.e., how effectively the features can be separated by a linear boundary. Our analysis reveals a strong linear correlation between k-NN performance and alignment score, as illustrated by an approximated line in the figure. In contrast, Linear Probing (see Fig. 10 in Supp.) shows a weaker correlation with alignment scores . Specifically, Pearson’s correlation between alignment accuracy (computed using our proposed metric) and ImageNet classification accuracy is 0.991 for k-NN classification and 0.847 for linear probing, highlighting that non-linear separability (i.e., clustering quality) matters more than linear-separability for image-text alignment.
@@
3 Part II: Learning Alignment between Unimodal Models
@@ 3.2.2 Zero-Shot Image-Text Retrieval
Cross-modal retrieval is more challenging than image recognition, as it requires complex scene understanding, including spatial relationships, context, background, activities, and more. Unlike models like CLIP that lack complex reasoning capabilities [40, 47, 1] as they train from scratch using large, noisy image-text datasets with short captions mentioning just object names, SAIL leverages a robust pretrained language encoder, trained for complex language understanding. With minimal alignment learning on a significantly smaller image-text dataset, SAIL achieves significant cross-modal understanding improvements.
@@
We further analyze the image-image cosine similarity on the MMVP benchmark [42], which consists of 150 image pairs. These pairs are selected to test subtle differences in orientation, perspective, quantity, color, and contextual details. Previous findings [42, 1] indicate that CLIP struggles with such distinctions, often assigning very high similarity scores even when condition varies, whereas DINOv2 effectively captures these nuances. As shown in Fig. 6, we found that SAIL’s cosine similarity distribution closely matches DINOv2’s, suggesting it retains DINOv2’s fine-grained visual acuity. By combining DINOv2’s visual precision with NV2’s linguistic depth, SAIL proves to be a powerful vision-language foundation model for nuanced visual discrimination.
@@ 3.2.3 Open-Vocabulary Semantic Segmentation
LAION400M CLIP [21] 1.2 2.4 15.8
LAION400M MaskCLIP [21] 6.9 8.9 30.1
LAION400M SCLIP [21] 7.1 13.1 60.3
@@
Table 3: Open-vocabulary semantic segmentation mIOU results compared with CLIP-based methods. All models use ViT-L/14 as the vision architecture.
@@ 3.2.4 Language-compatible Visual Representation
Figure 9: CLIP and LiT require training full models, making them resource-intensive. LiT and ShareLock freeze vision models during training, unable to yield language-compatible visual features crucial for MLLMs. CLIP and LiT achieves modality-specific representations and cross-modal alignment simultaneously, incapable for assessing alignment potential. Our framework meets all three requirements.
@@ 4 Related Work
Alignment between unimodal models.
Recent studies reveal that alignment can emerge within unimodal models even without explicitly aligning them with each other. Huh et al. [17] use mutual nearest-neighbor metrics to suggest that models across modalities align to a shared statistical reality. Maniparambil et al. [30] find that vision encoders exhibit high semantic similarity with language encoders using centered kernel distance. While these studies imply inherent alignment within unimodal models, they rely on proxy measurements without directly assessing cross-modal distance for individual image-text pairs. Moschella et al. [31] introduce relative representation, encoding samples based on similarity to a fixed set of anchors, ASIF [33] further extends this for cross-modal retrieval, requiring no training. Our approach directly measures the cross-modal distance for individual image-text pairs to quantitatively examine how modality-specific features impact alignment.
@@
of foundational VLMs like CLIP has remained underexplored. One line of work improves data efficiency by using improved captions from LLMs [10] or MLLMs [49], though computational demands remain high. Other methods combine pre-trained unimodal models to reduce data and compute needs. For instance, LiT [44] aligns a frozen vision model with a language model trained from scratch, which still requires substantial effort. Khan and Fu [18] explore parameter-efficient contrastive alignment by tuning layer norms and bias terms; ShareLock [36] further introduces a tunable head over the frozen language model to achieve alignment with fewer parameters. However, these methods have limited alignment performance and do not improve the base vision encoder, restricting the ability to transfer improved vision-language alignment capabilities to MLLMs. As shown in Fig. 9, our framework stands out by meeting three key points compared to other frameworks: efficient training, learning language-compatible visual features, and a more direct method for assessing alignment potential, consistent with how inference is done with such VLMs.
@@ 5 Conclusion
We propose a probing framework to evaluate cross-modal alignment between pretrained unimodal vision and language models and to explore how modality-specific features impact this alignment. We show that the clustering quality of SSL features is crucial for alignment and that advanced language models are essential for reasoning in vision-language tasks. We present SAIL, a framework that efficiently learns alignment with less data and compute demands. SAIL excels in zero-shot classification, retrieval, complex reasoning, and segmentation, outperforming CLIP trained on 400M image-text pairs with only 23M data. This study highlights the potential of efficient alignment strategies for advancing vision-language integration, aiming to accelerate research for resource-limited academic groups on topics such as architecture, losses, and data.
@@ Appendix A Reproducibility Statement
To ensure the reproducibility of our work, we are committed to making all training code, datasets, and model weights publicly available at Project Page. Detailed documentation will accompany the codebase to facilitate easy replication of our experiments. Hyperparameter settings, training configurations, and any preprocessing steps will also be thoroughly outlined. By providing these resources, we aim to promote transparency, enable future research, and support the broader community in building upon our work.
@@ Appendix C Additional Comparison with ShareLock
Experiments without enhanced caption
We provide additional results in Tab. 5 by training SAIL on CC12M raw dataset without long captions. We see that SAIL still significantly outperforms both ShareLock and LiT across all benchmarks, further validating the effectiveness of our method.
MSCOCO Flickr30k Winoground MMVP ImageNet 10 Classification
Data Model (DINOv2-B) I2T T2I I2T T2I T. I. G. 10 Avg. Top1. Avg.
[…]
CC12M raw ShareLock (NV2) 39.6 23.1 68.1 49.3 33.25 13 9.75 15.56 61.9 62.0
CC12M raw SAIL-B (NV2) 45.6 32.9 74.2 60.6 35.0 19.0 14.25 25.2 69.2 66.4
Table 5: Trained on CC12M raw captions.